OpenStudy (anonymous):

What polynomial identity should be used to prove that 25^2 = (20 + 5)2?

OpenStudy (solomonzelman):

\(25^2=(20+5)^2\) why wouldn't 20+5=25 qualify as a proof?

OpenStudy (solomonzelman):

25²=25² ....

OpenStudy (anonymous):

Difference of Cubes Difference of Squares Square of Binomial Sum of Cubes

OpenStudy (anonymous):

Those r the answer choices

OpenStudy (solomonzelman):

There is only 1 way to expand (20+5)², and that is, knowing that: \((a+b)^2=a^2+2ab+b^2\) Now, this is something that you would name using which of the options?

OpenStudy (anonymous):

square bionomial

OpenStudy (anonymous):

@SolomonZelman where'd ya go?

OpenStudy (solomonzelman):

yes

OpenStudy (solomonzelman):

Square of binomial is correct